OpenStudy (dtan5457):

What is the probability of getting a number of less than 3 on 6 out of 10 tosses of a fair die?

OpenStudy (dtan5457):

How would I set this up using binomial probability?

OpenStudy (dtan5457):

@butterflydreamer

OpenStudy (butterflydreamer):

i'm sorry but i can't help you :( I haven't studied binomial probability yet ....GOOD LUCK on this question though :))!

OpenStudy (dtan5457):

np thank you :)

OpenStudy (rational):

A die has numbers : \(\{1, 2, 3, 4, 5, 6\}\) We call it a success if we get less than 3 : \(\{1, 2\}\) So probability of success, \(p = 2/6 = 1/3\) Since we're rolling it 10 times, \(n=10\) \[P(X=6)~~ =~~ \binom{10}{6}(1/3)^6 (1-1/3)^{10-6}\] http://www.wolframalpha.com/input/?i=%5Cbinom%7B10%7D%7B6%7D%281%2F3%29%5E6+%281-1%2F3%29%5E%7B10-6%7D&dataset=

OpenStudy (kropot72):

\[\large 10C6 \times \frac{1}{3}^{6}\times\frac{2}{3}^{4}\]

OpenStudy (dtan5457):

Got it, thanks